OpenStudy (rizags):

HELP PLZ Each of the 50 states has two senators. In order for a bill to pass, it must have at least 50 votes. Suppose a bill passes with the minimum number of votes. Compute the probability that at least one Texan senator voted yes.

OpenStudy (anonymous):

does that mean compute the probability ?

OpenStudy (rizags):

yes

OpenStudy (rizags):

and the fact that he is texan has nothing to do with anything (just saying)

OpenStudy (rizags):

yes

OpenStudy (anonymous):

i think that maybe we pass to the compliment, what is the probability that no texas senator votes yes this is kind of a weird question, dons’t really relate to probability, more to combinatorics

OpenStudy (rizags):

yes i think it is more in the realm of combinatorics

OpenStudy (rizags):

im not sure how to interpret the statement "at least one"

OpenStudy (rizags):

is that the same as finding the probability that one senator voted yes, or that both voted yes

OpenStudy (anonymous):

the simple minded and probably wrong way to think about it is that the probability that any senator voted yes is \(\frac{1}{2}\) so the probability that at least one texas senator voted yes is \(\frac{3}{4}\) but i would not bet any money on this answer

OpenStudy (rizags):

wait one sec. I have another question, i must answer one of these two.

OpenStudy (anonymous):

i have to add that this question assumes something that is completely incorrect, namely that their votes are independent, which is simply a false assumption

OpenStudy (rizags):

Suppose that f and g are quadratic functions with f(8)=f(9)=100 and g(2)=g(4)= -20 and f(5)+g(0) is not equal to 0. Compute \[\frac{ f(12) + g(6) }{ f(5) + g(0) }\]

OpenStudy (rizags):

could you help me with this one then

OpenStudy (anonymous):

yeah i can do that the bone headed way for sure since apparently it doesn't matter what the actual function is we can make one up

OpenStudy (anonymous):

\[g(2)=g(4)=-20\] means that the function is symmetric about \(x=3\) so \(g(6)=g(0)\)

OpenStudy (rizags):

ok, makes sense, thanks. what can i do for f?

OpenStudy (rizags):

is this one symmetric about x=8.5?

OpenStudy (anonymous):

maybe the same thing? that one has \[f(8)=f(9)\] so that one is symmetric about \(x=8.5\) right

OpenStudy (anonymous):

that means that (i was hoping this would work) that \[f(5)=f(12)\] since \(8.5\) is right in the middle

OpenStudy (anonymous):

hence the numerator equals the denominator

OpenStudy (rizags):

OHHHHHHHHHH, so if i define f(5)=f(12)=x, and g(6)=g(0)=y, then i get the equation \[\frac{ x+y }{ x+y }=1\]

OpenStudy (rizags):

right?

OpenStudy (anonymous):

yup

OpenStudy (rizags):

THAAAANKs!!!!!